OpenStudy (anonymous):

one-ninth of polymer plastics sales are made in New England.If New England sales amount to $600,000, what are the total sales of the company?

jimthompson5910 (jim_thompson5910):

Let x = total sales (1/9)*(Total Sales) = (Sales in New England) (1/9)*(x) = (600,000) x = 9*(600,000) ... Multiply both sides by 9 to isolate x. x = 5,400,000 So the total sales is $5,400,000 (or 5.4 million dollars)